### v1

@@ Abstract
In this study, we identify and relax the assumption of data “poolability” in difference-in-differences (DID) estimation. Poolability, or the combination of observations from treated and control units into one dataset, is often not possible due to data privacy concerns. For instance, administrative health data stored in secure facilities is often not combinable across jurisdictions. We propose an innovative approach to estimate DID with unpoolable data: UN–DID. Our method incorporates adjustments for additional covariates, multiple groups, and staggered adoption. Without covariates, UN–DID and conventional DID give identical estimates of the average treatment effect on the treated (ATT). With covariates, we show mathematically and through simulations that UN–DID and conventional DID provide different, but equally informative, estimates of the ATT. An empirical example further underscores the utility of our methodology. The UN–DID method paves the way for more comprehensive analyses of policy impacts, even under data poolability constraints.
Keywords: difference-in-differences, treatment effects, siloed data, unpooled data, impact evaluation JEL Codes: C10, C12, C21, D04, I18, I38, Z18
@@ 1 Introduction
As economists, we are interested in evaluating the impact of a policy, both in attaining its intended objectives and in its unintended consequences. If the implementation or enforcement of a policy is randomized, we can easily estimate the average treatment effect on the treated (ATT) by comparing the means of the group that receives treatment and the group that does not. Under the more common non-randomized assignment to treatment, however, calculating the ATT is not straightforward due to selection bias. When policy changes are plausibly quasi-randomly assigned, difference-in-differences (DID) can be used to reasonably eliminate the selection bias and estimate the ATT.
@@
The set of identifying assumptions for estimating the ATT in the DID framework includes parallel trends, no anticipation, no staggered adoption, and homogeneous treatment effect (Abadie, 2005; De Chaisemartin and d’Haultfoeuille, 2020a; Callaway and Sant’Anna, 2021). In the past five years, developments in the DID literature have explored what happens when each assumption is violated and how to identify the ATT under each violation. The source of identifying variation in DID analyses is often a policy change that occurs in some political jurisdictions (e.g., states, countries) but not in others. Treated observations come from jurisdictions or units in which the policy change occurs, while control or untreated observations come from jurisdictions that do not experience that change (at least not at the same time).
The strong correlation between treatment assignment and jurisdiction implies another important assumption that has not been addressed in the DID literature: the data from the treatment group and the control group can be combined together for analyses. If the identifying variation is supplied by policy variation across countries and the data come from an international survey, for example, this assumption can be met. Observations from both treated and control countries are already pooled together in the same dataset. However, in settings where the data come from the same jurisdictional level as the policy variation, country-level administrative data, for example, data from treated countries may not be poolable with data from control countries. In this case, the DID assumption of poolable data is not met, and DID cannot be used to estimate the ATT of a policy that varies at the country level.
In recent years, the types and sources of available data have expanded and now include detailed genetic, income, and mobility data, among others. In some cases, privacy concerns and data regulations result in such data being “siloed,” that is, combining data across jurisdictions or with other data sources even within jurisdictions is restricted or prohibited (Li et al., 2022). The disaggregated data needed to estimate the ATT with the conventional DID regression cannot be shared between treated and untreated jurisdictions and combined together on a single server. Therefore, using current DID approaches to estimate the ATT for policies that vary by jurisdiction and where the data are siloed at that same jurisdictional level is difficult, if not infeasible.
One example of within-country siloed data is in Canada, where the universal health-care system is decentralized among 13 provinces and territories, each overseeing health-care policy and delivery. Each jurisdiction has a “single payer” health insurance plan, which generates population-based claims data that contain individual characteristics and health-care services funded by the public insurer. Combining the changing and varied health policies and structures across Canadian jurisdictions with the DID methodology provides an excellent opportunity to understand their impacts in order to improve health-care system performance and population health. However, data poolability constraints have limited most such analyses to using within-province controls, to comparing first differences in treated and control provinces or to using surveys or other pooled data sources.
Similar data poolability challenges are encountered beyond health care and internationally across countries. For example, certain country-level datasets, like Canada’s Vital Statistics Mortality database and the United States’ administrative tax records, are housed in secure facilities, thereby siloing them and hindering their poolability with similar datasets from other countries. The issue of federated data extends beyond health economics to various research domains, including medical research, epidemiology, environmental science, social science, data science, and policy analysis. In cases where a policy, like the legalization of marijuana in Canada, impacts a whole country, employing a different country as a control group may be ideal to evaluate the policy’s impacts. Under data poolability constraints, the full benefits of DID estimation cannot be realized. Researchers often estimate treatment effects using within-jurisdiction variation, which may be more prone to selection bias, or by employing interrupted time series analysis (Bernal et al., 2017), which can generate biased estimates if other changes coincide with the intervention of interest.
Several journal articles in data science, medical engineering, and computer science have explored how siloed data can pose a barrier to research (Li et al., 2022; Hallock et al., 2021; Wang and Alexander, 2020). The heterogeneity of data distributions between silos, implying that the data may not be independently or identically distributed, makes analysis with siloed data difficult (Li et al., 2022). Siloed data likely reduces the adoption of hospital information systems, which further restricts data diffusion (McCullough, 2008). The literature has addressed a need for access to and poolability of health datasets (Hallock et al., 2021; Miller and Tucker, 2014; Wang and Alexander, 2020; McCullough, 2008). However, no papers have explored how to use siloed data to conduct DID analysis.
In this paper, we introduce a DID estimator that estimates the ATT with unpoolable data (UN) from the treated and control groups, which we call UN–DID. Because researchers commonly rely on regression-based tools for estimating treatment effects, the primary goal of this paper is to introduce the UN–DID estimator as a regression-based tool for this purpose. Through analytical proofs and a simulation study, we demonstrate that UN–DID can estimate relevant ATTs and their associated standard errors. The estimated ATTs from the UN–DID estimator are unbiased across various data-generating processes and converge to a common value with large sample sizes. This holds when we include both time-invariant and time-varying covariates. We also showcase a practical application of our methodology using two empirical examples.
@@ 2 Theoretical Framework
Let us assume that we have data for two groups, the treatment group and the control group, and data for several calendar years, $t=1,2,.....,T$. We also assume that the treatment was implemented at year $k$. Groups that received the treatment are called the treated group and groups that did not receive the treatment are called the control group. All calendar years before $k$ ($t<k$) are pooled together as the pre-intervention period. Similarly, all calendar years from $k$ onward ($t\geq k$) are pooled together as the post-intervention period. It is common practice to estimate DID using micro-level data of this nature, where numerous observations from individual subjects are available within each of the relevant cells.
The basic idea of DID estimation is to compare the difference in outcomes before and after the treatment between groups that received the treatment and groups that did not (Bertrand et al., 2004). The difference in outcomes between the treatment group and the control group before treatment is used to impute the unobserved counterfactual of the treated group. Card and Krueger (1993) first used the conventional DID to estimate the effect of an increase in minimum wage on unemployment in the American state of Pennsylvania. The traditional way of estimating the ATT involves running the following conventional regression:
@@
Here, $D^{S}_{i}$ is a dummy variable that takes on a value of 1 if individual $i$ is in the treatment group, and 0 otherwise. $P^{S}_{t}$ is a dummy variable that takes on a value of 1 if the data is for post-intervention period, and 0 otherwise. $S$ implies that we are combining the dataset from both the treated group $T$ and the control group $C$ into a common dataset ($S=T+C$). Under very strong assumptions, $\hat{\beta_{3}}$ (the coefficient of the interaction term between $D^{S}_{i}$ and $P^{S}_{t}$) identifies the ATT (Roth et al., 2022). To simplify notations, we will not use the $S$ superscripts to denote a combined dataset for the rest of the paper.
@@
$\begin{split}&\biggr{[}E[Y_{i}(0)|D_{i}=1,t=1]-E[Y_{i}(0)|D_{i}=1,t=0]\biggr{]}\\ &=\ \biggr{[}E[Y_{i}(0)|D_{i}=0,t=1]-E[Y_{i}(0)|D_{i}=0,t=0]\biggr{]}.\end{split}$ (2)
@@
In recent years, much work has been done in the DID literature to investigate what occurs when each of the aforementioned assumptions are violated. In this paper, we introduce a new assumption that has been implied in previous literature but has not been explicitly explored or discussed in the context of DID.
Assumption 6: Data is Poolable
Data from the treatment and control groups are available on a single server and can be combined together for analyses. Let $T$ and $C$ be two data silos, where $T$ is the treated silo and $C$ is the untreated silo. $Y_{T}$ and $Y_{C}$ are two matrices containing data for the outcome variable for silo $T$ and $C$, respectively, and $N_{T}$ and $N_{C}$ the total number of observations in the matrices, respectively.
@@
Under the poolability assumption, the two matrices can be stacked together into a common matrix Y, which is not feasible when there are legal restrictions preventing them from being stacked together (or data are unpooled):
@@
For situations where Assumption 6 is violated, we introduce the UN–DID approach, which can estimate the ATT in scenarios where the data is not poolable. With unpoolable data, when we access the data from the treated silo, we have no knowledge of the data from the control group, $Y_{C}$. Similarly, once we have access to the data from the untreated silo, we have no information about the data from the treated silo, or $Y_{T}$. We know only the treatment status of the individuals in each silo and whether the data are from the pre- or post-intervention period. The conventional method becomes impractical in situations when data are segregated, or is “siloed,” because datasets cannot be combined for conventional regression analysis.
In this case, we can visit each silo and run regressions shown in equations (6) and (7) for the treated and untreated silos, respectively. In the simplest case, we assume that there are only two silos: the treated silo ($T$) and the untreated silo ($C$). We also assume that there are no covariates. Here, $Y^{T}_{i,t}$ is the outcome for an individual from the silo that is treated at time $t$ and $Y^{C}_{i,t}$ is the outcome of an individual from the silo that is untreated at time $t$. $post_{t}^{T}$ is a dummy variable that takes on a value of 1 when the treated observation is in the post-intervention period, and 0 otherwise. Similarly, $post_{t}^{C}$ is a dummy variable that takes on a value of 1 when the untreated observation is in the post-intervention period, and 0 otherwise. $pre_{t}^{T}$ is a dummy variable that takes on a value of 1 if the treated observation is in the pre-intervention period, hence $pre_{t}^{T}=1-post_{t}^{T}$. Similarly, $pre_{t}^{C}$ is a dummy variable that takes on a value of 1 if the untreated observation is in the pre-intervention period, hence $pre_{t}^{C}=1-post_{t}^{C}$. Note that the regressions in equations (6) and (7) are done without a constant. Therefore, none of the variables are dropped because of multicollinearity.
@@
These regressions are silo-specific. Because data are siloed, the treated regression does not contain any data from the untreated silo and the untreated regression does not contain any data from the treated silo. Here, $(\widehat{\lambda_{2}^{T}}-\widehat{\lambda_{1}^{T}})-(\widehat{\lambda_{2}^{C}}-\widehat{\lambda_{1}^{C}})$ is the estimate of the ATT, as shown in Equation (8). The associated standard error of the ATT is estimated as the square root of the sum of the standard errors used to estimate the ATT estimated from the above UN–DID regressions. This is shown in Equation (9).
@@
$\widehat{SE}(\widehat{ATT})=\sqrt{\widehat{SE}(\widehat{\lambda_{1}^{T}})^{2}+\widehat{SE}(\widehat{\lambda_{2}^{T}})^{2}+\widehat{SE}(\widehat{\lambda_{1}^{C}})^{2}+\widehat{SE}(\widehat{\lambda_{2}^{C}})^{2}}.$ (9)

### v2

@@ Abstract
Difference-in-differences (DID) is commonly used to estimate treatment effects but is infeasible in settings where data are unpoolable due to privacy concerns or legal restrictions on data sharing, particularly across jurisdictions. In this study, we identify and relax the assumption of data poolability in DID estimation. We propose an innovative approach to estimate DID with unpoolable data (UN-DID) which can accommodate covariates, multiple groups, and staggered adoption. Through analytical proofs and Monte Carlo simulations, we show that UN-DID and conventional DID estimates of the average treatment effect and standard errors are equal and unbiased in settings without covariates. With covariates, both methods produce estimates that are unbiased, equivalent, and converge to the true value. The estimates differ slightly but the statistical inference and substantive conclusions remain the same. Two empirical examples with real-world data further underscore UN-DID’s utility. The UN-DID method allows the estimation of cross-jurisdictional treatment effects with unpoolable data, enabling better counterfactuals to be used and new research questions to be answered.
Keywords: difference-in-differences, treatment effects, siloed data, unpooled data, impact evaluation
JEL Codes: C10, C12, C21, D04, I18, I38, Z18
@@ 1 Introduction
Difference-in-differences (DID) is a widely used method to evaluate the impacts of policies and interventions. It is used to estimate both intended treatment effects and unintended consequences. The source of identifying variation in DID analyses is often a policy change that occurs in some political jurisdictions (e.g., states, countries) but not in others. Treated observations come from jurisdictions or units in which the intervention occurs, while control or untreated observations come from jurisdictions that do not experience that change (at least not at the same time). DID can be used to estimate both the intent-to-treat (ITT) effect and the average treatment effect on the treated (ATT). In the absence of randomized treatment assignment, a widely documented set of assumptions support the unbiased estimation of these effects. These assumptions include parallel pre-intervention trends, no anticipation, no staggered adoption, homogeneous treatment effects, and a single isolated treatment (Roth et al., 2022; Abadie, 2005; De Chaisemartin and d’Haultfoeuille, 2020a; Callaway and Sant’Anna, 2021; De Chaisemartin and D’Haultfœuille, 2020b). The strong correlation between treatment assignment and jurisdiction implies an additional assumption underlying the use of DID estimation that has not yet been made explicit: that data from the treatment group and the control group can be combined together for analyses.
This paper develops the DID for unpoolable data (UN-DID) method, enabling the estimation of treatment effects using DID when the data pooling assumption is violated. This occurs when data come from the same jurisdictional level as the policy variation, and may not be combined on the same computer or server, usually due to privacy concerns or legal restrictions on data sharing. The data are therefore siloed at the level of treatment, and traditional DID methods cannot be used to estimate the ATT. Examples include health care data from different insurers (e.g., Canadian provinces (Health Data Research Network Canada, 2021), private and public insurers, all-payer claims databases across states (Kranz et al., 2023)), restricted-use micro-data held in secure environments (e.g., vital statistics, income tax, business data held in Statistics Canada’s or the US Census Bureau’s Research Data Centres (Canadian Research Data Centre Network, 2024; United States Census Bureau, 2024)), and electronic health records federated across countries (e.g., the European Health Data Space (European Commission, 2024)). Sometimes data are considered unpoolable because the datasets are not sufficiently harmonized. We do not address that issue here - UN-DID requires that data that are physically unpoolable be sufficiently harmonized (either de facto or after cross-jurisdictional collaboration). That said, UN-DID is able to accommodate datasets from different jurisdictions that measure the same underlying concepts, but code or categorize variables slightly differently (e.g., income in different bins). UN-DID enables the estimation of treatment effects when policies and interventions vary at the level of data silos (county, state, country, etc.), opening opportunities to estimate less-biased treatment effects using better controls and to answer new research questions.
When the ability to pool data across treatment and control jurisdictions is limited, the choice of appropriate counterfactuals is necessarily constrained. Researchers may resort to estimating first-differences (Fischer et al., 2021) or using interrupted time series (Tamblyn, 2001; Bernal et al., 2017; Farin et al., 2024), which only provide unbiased treatment effect estimates if no other changes coincide with the intervention of interest. More convincing treatment effect estimates are obtained leveraging within-jurisdiction controls (Rudoler et al., 2023; Strumpf et al., 2017; Li et al., 2014; Girma and Görg, 2007) or individual fixed effects (Dumont et al., 2008), both of which may still suffer from selection on unobservables. The impacts of national policies that take effect everywhere at the same time can sometimes be identified using pre-policy variation (Finkelstein, 2007; Duflo, 2001), but in other cases, including nationwide cannabis legalization in Canada in 2018, rigorous policy evaluations using external controls remain extremely limited or non-existent.
If policy designs permit the use of RD or IV to estimate local average treatment effects (Almond et al., 2010; Hutcheon et al., 2020), the identifying variation may be convincing, but the generalizability of the estimate narrow by definition. Meta-analysis can be used with unpoolable data, and is an effective way to leverage sample size by combining silo-specific treatment effect estimates into a single parameter (Xiong et al., 2023). For example, the Canadian Network for Observational Drug Effect Studies (CNODES) routinely employs meta-analytical techniques to more comprehensively examine drug safety using siloed provincial or national data (Suissa et al., 2012; Filion et al., 2014; Durand et al., 2021). However, meta-analysis cannot be used to estimate treatment effects of an intervention in one jurisdiction using another unexposed jurisdiction as a control. UN-DID will allow researchers to estimate treatment effects using better counterfactuals, leveraging data from non-treated jurisdictions and limiting selection bias while maintaining generalizability.
Data poolability limitations can also constrain which research questions get asked, specifically which outcomes are examined in evaluations of policy interventions. National and international surveys include multiple states/provinces or countries, so DID models are easily estimated with treated and control units in the same dataset (Baker et al., 2008; Charters et al., 2013; Dube, 2019). The range of outcomes available in a survey, however, are necessarily limited with respect to content, detail, and method of reporting (self-report vs. administrative data). In the absence of a method like UN-DID to allow treatment effect estimation across separate datasets, the questions that have been asked and answered have been a function of the (pooled) datasets that are available. This constraint has become even more problematic in recent years, precisely because large, detailed, and high-quality data are increasingly available for research. All-payer health insurance claims, electronic medical records, detailed individual- and firm-level tax filings, genetic data, and immigration records offer comprehensiveness, accuracy and detail that allow researchers to ask and answer new questions. However, these data are generally accessible only in secure research data centers and are not shareable or poolable across jurisdictions. By enabling treatment effect estimation using siloed data, UN-DID will open up opportunities to answer novel research questions and exploit new sources of identifying variation using separate datasets across jurisdictions.
The UN-DID method allows the estimation of treatment effects with unpoolable data. This has important implications for research across the fields of applied economics, and across a range of disciplines including public policy, epidemiology, health research, environmental science, political science, sociology, and others. With UN-DID, new research questions can be answered and better counterfactuals can be used to reduce bias when estimating treatment effects. While researchers with treatment and control observations in separate datasets could already calculate the double-difference “by hand”, our formal results justify this approach and extend beyond the 2$\times$2 case. Furthermore, UN-DID enables the estimation of appropriate standard errors and therefore proper inference. We are developing UN-DID software packages and a user guide that will follow as a separate publication. These will enable UN-DID to be implemented across secure data silos, even when local data analysts are not skilled econometricians.
The intuition behind UN-DID is based on non-parametric DID estimation where conditional mean outcomes are estimated for each of four treatment*timing groups: pre-intervention treatment, pre-intervention control, post-intervention treatment and post-intervention control. The post-pre first difference can then be calculated for each treatment and control group, and then the DID estimate is simply the difference between those first differences. UN-DID estimates first-differences within jurisdictional silos and provides a method to combine these estimates into a treatment effect estimate and to test hypotheses. We provide formal results which show that this method accurately estimates the ATT. UN-DID also estimates the parameters necessary to evaluate parallel pre-intervention trends.
In this paper, we begin by reviewing the conventional DID identifying assumptions and formally introduce the poolability assumption, that data can be combined across treatment and control groups. We start with a one treatment, one control, and several time periods (2$\times$×T) setting. We propose a method for estimating the ATT when the poolability assumption is violated, which we call UN-DID (DID for unpoolable data). Because researchers commonly rely on regression-based tools for estimating treatment effects, we introduce the UN-DID method as a regression-based tool for this purpose. Through analytical proofs, we first demonstrate the equivalence of estimates between conventional DID and UN-DID without covariates. We then relax the strong parallel trends assumption and add the assumption of common causal covariates (each covariate has a common slope across silos). In this setting, we show the equivalence of conventional DID and UN-DID with time-invariant covariates, and in larger samples with time-varying covariates. Lastly, we turn to the many group many time period (G×$\times$×T) setting and relax the no staggered adoption assumption. We show how the UN-DID method works in the staggered adoption setting in the spirit of Callaway and Sant’Anna (2021). Throughout, we also demonstrate the equivalence of conventional standard errors between conventional DID and UN-DID. Cluster robust inference is possible as described in work in progress by the authors of this paper.
With analytical results in hand, we conduct Monte Carlo experiments with simulated panel data designed to mimic populations in Canada’s two largest provinces over 10 years (2$\times$10). We consider data generating processes (DGPs) with and without treatment effects and with no covariates, with a time-fixed and with a time-varying covariate. We consider equal and unequal sample sizes between the two provinces, and vary the sample size from 500 to 50,000 observations. To assess unbiasedness, we plot kernel densities of UN-DID and conventional DID ATT estimates against the known ATT value from the DGP. To assess the asymptotic performance of UN-DID and the conventional DID, we compare estimates with the true values of the ATT and standard errors against to the known true values as the sample size increases. Specifically, we calculate the mean squared error between the ATTs and the SEs for each pair: UN-DID vs. truth, convential DID vs. truth, and UN-DID vs. conventional DID. Lastly, to assess UN-DID’s performance and to highlight its utility, we include two empirical examples: one with common treatment timing and one with staggered adoption. Here, real poolable data are treated as if they were unpoolable across treated and control groups, allowing us to compare the estimated ATTs and SEs from UN-DID and conventional DID with real-world data.
In the simulation study, we demonstrate that UN-DID and conventional DID are both unbiased and asymptotically consistent without covariates. With a time-invariant covariate, the two ATTs are unbiased and converge to the true value; the SEs are equivalent but not numerically equal. With a time-varying covariate, the UN-DID and conventional DID ATTs and SEs are unbiased, equivalent, and converge to the true value. The above results hold with both equal and unequal sample sizes across silos. These results hold in our two empirical examples using real-world data. Without covariates, UN-DID’s ATT and SE estimates are exactly the same as conventional DID. With covariates, the estimates differ slightly but the statistical inference and substantive conclusions remain the same. As the number of covariates and/or data silos increases, the more flexible UN-DID model may differ more from conventional DID which constrains covariates to have the same slopes across silos. These results reinforce the theoretical and simulation study findings, that the UN-DID procedure estimates either identical or very similar treatment effects compared to the conventional method.
This paper contributes to the dynamic literature on DID methods. In the past five years, an important series of papers have looked under the hood of the conventional DID estimator, exploring what is actually being estimated and what can go wrong when each of the identifying assumptions is violated (Roth et al., 2022; Abadie, 2005; De Chaisemartin and d’Haultfoeuille, 2020a; Callaway and Sant’Anna, 2021; De Chaisemartin and D’Haultfœuille, 2020b; Goodman-Bacon, 2021). They have developed methods to identify the ATT in each case, changing the way that DID analyses are conducted. We highlight the heretofore implicit assumption of data poolabilty and provide a solution to estimate the ATT when this assumption does not hold. We assess UN-DID’s performance and demonstrate its equivalence to conventional DID.
We evidently aim to contribute to the large literature that uses DID to estimate treatment effects of policy changes and other interventions, both in economics and across a wide range of disciplines. In this paper, we apply UN-DID to two published policy evaluations that used conventional DID methods (Sabia et al., 2017; Conley and Taber, 2011). Our primary goal is to compare the estimates from the two methods, with a secondary goal of highlighting how UN-DID could be used in empirical analyses. The forthcoming software package user guide will use UN-DID to evaluate a policy change using real-world siloed data, where a conventional DID analysis is impossible and no ATT estimates exist for comparison.
The paper proceeds as follows: Section 2 presents the theoretical framework and the identifying assumptions underlying conventional DID. Section 3 demonstrates the equivalence of the ATT estimated with UN-DID and with conventional DID when the strong parallel trends assumption holds, describes the basic Monte Carlo design, and presents simulation results for the cases without covariates. Section 4 demonstrates the equivalence UN-DID and conventional DID under the weaker assumption of conditional parallel trends and of common causal covariates, expands the Monte Carlo design to include time-fixed and time-varying covariates, and presents results. Section 5 presents UN-DID with staggered adoption. Section 6 describes a simplified method of aggregating estimates from multiple silos. Section 7 compares UN-DID and conventional DID ATT estimates for two empirical examples using real-world data and Section 8 concludes.
@@ 2 Theoretical Framework
Let us assume that we have data for two groups, the treatment group which received the intervention or treatment and the control group which did not. We also assume data for several calendar years, $t=1,2,.....,R$ and that the treatment was implemented at year $k$. All calendar years before $k$ ($t<k$) are pooled together as the pre-intervention period. Similarly, all calendar years from $k$ onward ($t\geq k$) are pooled together as the post-intervention period. It is common practice to estimate DID using micro-level data, where numerous observations from individual subjects are available within each of the relevant cells.
The basic idea of DID estimation is to compare the difference in outcomes before and after the treatment between groups that received the treatment and groups that did not (Bertrand et al., 2004). The change in outcomes for the control group is used as the unobserved counterfactual of the treated group. Card and Krueger (1993) first used the conventional DID to estimate the effect of an increase in minimum wage on unemployment in the American state of Pennsylvania. The conventional way of estimating the ATT uses the following regression:
@@
Here, $D^{S}_{i}$ is a dummy variable that takes on a value of 1 if individual $i$ is in the treatment group, and 0 otherwise. $P^{S}_{t}$ is a dummy variable that takes on a value of 1 if the observation is in the post-intervention period, and 0 otherwise. $S$ implies that we are combining the dataset from both the treated group $T$ and the control group $C$ into a common dataset ($S=T+C$). Under very strong assumptions, $\hat{\beta_{3}}$ (the coefficient of the interaction term between $D^{S}_{i}$ and $P^{S}_{t}$) identifies the ATT (Roth et al., 2022). To simplify notations, we will not use the $S$ superscripts to denote a combined dataset for the rest of the paper.
@@
$\begin{split}&\biggr{[}E[Y_{i}(0)|D_{i}=1,P_{t}=1]-E[Y_{i}(0)|D_{i}=1,P_{t}=0]\biggr{]}\\ &=\ \biggr{[}E[Y_{i}(0)|D_{i}=0,P_{t}=1]-E[Y_{i}(0)|D_{i}=0,P_{t}=0]\biggr{]}.\end{split}$ (2)
@@
The papers cited above have contributed to a rich recent literature investigating what occurs when each of the aforementioned assumptions are violated. In this paper, we introduce a new assumption that has been implied in previous literature but has not been explicitly explored or discussed in the context of DID.
Assumption 6: Data are Poolable
Data from the treatment and control groups are available on a single server and can be combined together for analyses. $Y_{T}$ and $Y_{C}$ are two matrices containing the outcome variable for group $T$ and $C$, respectively, and $N_{T}$ and $N_{C}$ the total number of observations in the matrices, respectively.
@@
Under the poolability assumption, the two matrices can be stacked together into a common matrix Y, which is not feasible when there are legal restrictions preventing them from being stacked together (data are siloed or unpoolable):
@@
For settings where Assumption 6 is violated and the data are not poolable, we introduce the UN-DID method to estimate the ATT. In such cases, when we access the data from the treated silo, we have no knowledge of the data from the control silo, $Y_{C}$. Similarly, once we have access to the data from the untreated silo, we have no information about the data from the treated silo, or $Y_{T}$. The conventional DID method becomes impractical if siloed datasets cannot be combined for regression analysis.
In this case, we can visit each silo (physically or virtually) and run regressions shown in equations (6) and (7) for the treated and untreated silos, respectively. In the simplest case, we assume that there are only two silos: the treated silo ($T$) and the untreated silo ($C$). We also assume that there are no covariates. Here, $Y^{T}_{i,t}$ is the outcome for an individual from the silo that is treated at time $t$ and $Y^{C}_{i,t}$ is the outcome of an individual from the silo that is untreated at time $t$. $post_{t}^{T}$ is a dummy variable that takes on a value of 1 when the treated observation is in the post-intervention period, and 0 otherwise. Similarly, $post_{t}^{C}$ is a dummy variable that takes on a value of 1 when the untreated observation is in the post-intervention period, and 0 otherwise. $pre_{t}^{T}$ is a dummy variable that takes on a value of 1 if the treated observation is in the pre-intervention period, hence $pre_{t}^{T}=1-post_{t}^{T}$. Similarly, $pre_{t}^{C}$ is a dummy variable that takes on a value of 1 if the untreated observation is in the pre-intervention period, hence $pre_{t}^{C}=1-post_{t}^{C}$. Note that the regressions in equations (6) and (7) do not include a constant. Therefore, none of the variables are dropped because of multicollinearity.
@@
These regressions are silo-specific. Because data are siloed, the treated regression does not contain any data from the untreated silo and the untreated regression does not contain any data from the treated silo. Here, $(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})$ is the estimate of the ATT, as shown in Equation (8). The associated standard error of the ATT is estimated as the square root of the sum of the variances used to estimate the ATT estimated from the above UN-DID regressions minus twice the covariances between the relevant coefficients. This is shown in Equation (9).
@@
$\small\widehat{SE}(\widehat{ATT})=\sqrt{\widehat{SE}(\widehat{\lambda}_{1}^{T})^{2}+\widehat{SE}(\widehat{\lambda_{2}}^{T})^{2}+\widehat{SE}(\widehat{\lambda}_{1}^{C})^{2}+\widehat{SE}(\widehat{\lambda}_{2}^{C})^{2}-2\widehat{Cov}(\widehat{\lambda}_{1}^{T},\widehat{\lambda}_{2}^{T})-2\widehat{Cov}(\widehat{\lambda}_{1}^{C},\widehat{\lambda}_{2}^{C})}.$ (9)